### v1

@@ 6 Policy Iteration for Pareto-Optimal Policies
SSE policies do not always exist in general-sum stochastic games. Therefore, it is desired that an algorithm converges to an SSE policy if it exists, and to a policy that is “reasonable” provided an SSE policy does not exist. First, we introduce the notion of Pareto-optimality as the reasonable target. A policy that approximates a Pareto-optimal value function with an arbitrary precision always exists and it admits an SSE policy if an SSE policy exists. Then, we derive a policy improvement theorem for SSGs. Based on the policy improvement theorem for SSGs, we design an algorithm that monotonically improves the policy toward Pareto-optimal policies.
@@ 6.1 Pareto-Optimal Policy
There always exists $\mathcal{PV}\neq\emptyset$. Moreover, $\mathcal{PV}\subseteq\partial\mathcal{V}$. A PO policy exists if and only if $\mathcal{PV}\cap\mathcal{V}\neq\emptyset$. For any PO value function $v^{*}\in\mathcal{PV}$, there always exists a policy whose value function is arbitrarily close to $v^{*}$. Meanwhile, there is no stationary policy that realizes a value function better than a PO value function. Moreover, an arbitrary PO policy is an SSE policy if there exist SSE policies, as shown in Proposition 6.2. Therefore, PO value functions are a reasonable alternative to the SE value function. The proof is presented in Appendix F.
@@ 6.2 Policy Improvement Theorem
Lemma 6.3 is an extension of the policy improvement theorem Sutton & Barto (2018) in single-agent MDPs. With this theorem, we can confirm if a policy $f$ is better than a baseline $f^{\prime}$ without computing the value function for $f$. The proof is shown in Appendix G.
Lemma 6.3.
@@
  • (a)
    $Q_{A}^{f^{\prime}{\dagger}}(\cdot,f)\succcurlyeq V_{A}^{f^{\prime}{\dagger}}\implies V_{A}^{f{\dagger}}\succcurlyeq Q_{A}^{f^{\prime}{\dagger}}(\cdot,f)\succcurlyeq V_{A}^{f^{\prime}{\dagger}}$;
  • (b)
    $Q_{A}^{f^{\prime}{\dagger}}(\cdot,f)\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f^{\prime}{\dagger}}\iff V_{A}^{f{\dagger}}\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f^{\prime}{\dagger}}$;
  • (c)
    $Q_{A}^{f^{\prime}{\dagger}}(\cdot,f)\succ V_{A}^{f^{\prime}{\dagger}}\iff\\ V_{A}^{f{\dagger}}-V_{A}^{f^{\prime}{\dagger}}(s)\succ\gamma_{A}\mathbb{E}_{s^{\prime}}\left[V_{A}^{f{\dagger}}(s^{\prime})-V_{A}^{f^{\prime}{\dagger}}(s^{\prime})\right]$.
A necessary condition for PO policies is provided below. It shows that any improvement by Lemma 6.3 (a) is not possible from PO policies. This proof is in Appendix H.
Lemma 6.4 (Necessary condition for PO policies).
@@
The following result shows a sufficient condition for PO policies. This proof is presented in Appendix I.
Lemma 6.5 (Sufficient condition for PO policies).
@@
  • i)
    $Q_{A}^{f^{\prime}{\dagger}}(s,f)<V_{A}^{f^{\prime}{\dagger}}(s)-\dfrac{\gamma_{A}}{1-\gamma_{A}}\delta(f,f^{\prime})$, $\exists s\in\mathcal{S}$, where $\delta(f,f^{\prime})=\max\limits_{s\in\mathcal{S}}\mathbb{E}_{s^{\prime}}\left[Q_{A}^{f^{\prime}{\dagger}}(s^{\prime},f)-V_{A}^{f^{\prime}{\dagger}}(s^{\prime})\right]$ and the expectation is considered for $s^{\prime}\sim p(s^{\prime}|s,f(s),R_{B}^{*}(s,f))$;
  • ii)
    $Q_{A}^{f^{\prime}{\dagger}}(\cdot,f)\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f^{\prime}{\dagger}}$.
@@ 6.3 Policy Iteration for PO Policies
We propose an approach that monotonically improves a policy toward PO policies and is guaranteed to converge.
PO value functions are not always unique. To guide the algorithm, we introduce scalarization. Given a stationary policy $f\in\mathcal{W}_{A}$, we maximize a Pareto-compliant scalarization $\mathcal{L}$ of the state value on $f$, namely, $\mathcal{L}[V_{A}^{f{\dagger}}]$, where the Pareto-compliant scalarization555 For example, a weighted sum scalarization $\mathcal{L}[v]=\sum_{s\in\mathcal{S}}\alpha_{s}v(s)$ is a Pareto-compliant scalarization for $\boldsymbol{\alpha}:=\{\alpha_{s}\in\mathbb{R}_{>0}\}_{s\in\mathcal{S}}$ such that $\sum_{s\in\mathcal{S}}\alpha_{s}=1$. If we choose $\alpha$ such that $\alpha_{s}=\rho(s)$ for all $s\in\mathcal{S}$, then $\mathcal{L}[v]$ coincides with the expected return under the initial state distribution $\rho$. $\mathcal{L}:\mathrm{cl}\mathcal{V}\to\mathbb{R}$ is a continuous function such that $v\succ v^{\prime}\implies\mathcal{L}[v]>\mathcal{L}[v^{\prime}]$. Then, the maximum of the scalarized value in $\mathrm{cl}\mathcal{V}$ is obtained by a PO value function, namely, $\operatornamewithlimits{argmax}_{v\in\mathrm{cl}\mathcal{V}}\mathcal{L}\left[v\right]\subseteq\mathcal{PV}$.
@@
Selecting the next policy $f_{t+1}$ from $\mathcal{W}_{\succcurlyeq}(f_{t})$ guarantees $\mathcal{L}[V_{A}^{f_{t+1}{\dagger}}]\geq\mathcal{L}[V_{A}^{f_{t}{\dagger}}]$ in light of Lemma 6.3 (a). Ideally,666 In practice, $\operatornamewithlimits{argmax}$ is not necessary, and sometimes it does not exist or is not computable. An alternative approach is to select a policy $f_{t+1}$ from a set $\mathcal{W}_{1-\epsilon}(f_{t})$ of policies that are $\epsilon$-optimal in terms of $\mathcal{L}$, such as $\mathcal{W}_{1-\epsilon}(f_{t}):=\bigg{\{}f\in\mathcal{W}_{\succcurlyeq}(f_{t})\mid\mathcal{L}\left[Q_{A}^{f_{t}{\dagger}}(\cdot,f)\right]-\mathcal{L}\left[V_{A}^{f_{t}{\dagger}}\right]\\ \geq(1-\epsilon)\sup_{f\in\mathcal{W}_{\succcurlyeq}(f_{t})}\bigg{(}\mathcal{L}\left[Q_{A}^{f_{t}{\dagger}}(\cdot,f)\right]-\mathcal{L}\left[V_{A}^{f_{t}{\dagger}}\right]\bigg{)}\bigg{\}}$ (9) for some $\epsilon\in[0,1)$. Notably, $\mathcal{W}(f_{t})=\mathcal{W}_{1}(f_{t})$. Therefore, the update in Equation 10 is a special case of Equation 9. If it is still intractable, selecting $f_{t+1}\in\mathcal{W}_{\succcurlyeq}(f_{t})$ s.t. $Q_{A}^{f_{t}{\dagger}}(\cdot,f_{t+1})\succ V_{A}^{f_{t}{\dagger}}$ ensures that (a) and (b) in Theorem 6.6 hold. the proposed algorithm selects $f_{t+1}$ from the set of policies that maximizes the scalarized value, namely,
@@
Theorem 6.6 guarantees that the proposed algorithm always converges and that the limit satisfies the necessary condition for PO policies (i.e., $v_{\infty}\in\partial\mathcal{V}$). The proof is shown in Appendix J.
@@
Let $f_{0}\in\mathcal{W}_{A}$ be an arbitrary initial policy and a policy sequence $\{f_{t}\in\mathcal{W}(f_{t-1})\}_{t=1}^{\infty}$ be obtained by Equation 9. Let $\{v_{t}=V_{A}^{f_{t}{\dagger}}\}_{t=0}^{\infty}$ be the corresponding value functions. Then, the following statements hold:
  • a)
@@
  • c)
    $v_{\infty}\in\partial\mathcal{V}$ and for all $v\in\mathcal{V}$,
    $\min_{s\in\mathcal{S}}v(s)-v_{\infty}(s)\leq\gamma_{A}\left(\max_{s\in\mathcal{S}}v(s)-v_{\infty}(s)\right)$;
  • d)
@@
As shown in the proof (Appendix J), the algorithm ceases to improve if and only if $Q_{A}^{f_{t}{\dagger}}(\cdot,f)\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f_{t}{\dagger}}$ for all $f\in\mathcal{W}_{\succcurlyeq}(f_{t})$. The algorithm terminates when it holds and returns $f_{t}$. The whole proposed algorithm summarizing the above is shown in Algorithm 1.
Algorithm 1 Policy Iteration for PO Policies (PIPOP)
@@
4:     Compute $\mathcal{W}_{\succcurlyeq}(f_{t})$ in Equation 8
@@
8:        Compute $\mathcal{W}(f_{t})$ in Equation 10 or Equation 9
@@
The convergence guarantee and the monotone improvement of Algorithm 1 (Statements (a) and (b) in Theorem 6.6) show the advantages of the proposed approach. However, compared with the policy iteration approach in single-agent MDPs, where the convergence toward the optimal policy is guaranteed, the fixed point of Algorithm 1 is not assured to be PO value functions. Rather, Theorem 6.6 (c) guarantees that the limit is at the boundary $\partial\mathcal{V}$, including all the PO value functions. In a special case of $\gamma_{A}=0$, the limit is guaranteed to be a PO value function (Theorem 6.6 (d)).
Furthermore, there exists an algorithmic difference to the policy iteration for the single-agent MDP. In single-agent MDPs, a state-wise maximization of the action value function of $f_{t}$, $a_{s}\in\operatornamewithlimits{argmax}_{a\in\mathcal{A}}Q^{f_{t}}(s,a)$ is performed to construct the next (deterministic) policy. This is invalid in SSGs because such a policy does not necessarily satisfy the condition of Lemma 6.3 (a) and may not improve the values. Moreover, the optimal policy is not necessarily a deterministic one. Because of these differences, Equation 10 cannot be simplified as the state-wise maximization. This can be a limitation from a practical viewpoint when Equation 10 (or Equation 9) is intractable.
In practice, it is helpful to restart the policy iteration process from other initial policies to obtain a better policy unless the obtained policy satisfies the sufficient condition Lemma 6.5. In this situation, one can randomly pick an initial policy or select a policy $f\in\mathcal{W}(f_{t-1})$ that is not yet explored. Because $f\in\mathcal{W}(f_{t-1})$ is already improved from the previous initial policy, we can expect such an initialization to be more efficient. The detail of the proposed approach with the restart strategy is provided in Appendix K.
@@ 7 Limitation and Future Work
This study proposed a novel algorithm with a convergence guarantee in general-sum SSGs. We introduced the notion of the Pareto-optimal value function and developed an algorithm that monotonically improves a policy toward the Pareto-optimal value functions. While existing methods have little guarantee, especially in games where the SSEs do not exist, our proposed approach can monotonically improve the leader’s performance in general, which is of great importance in practice. A potential application of the proposed approach is the intervention to an adaptive or learning follower agent as described in Appendix M, where the follower is not usually myopic; hence the existence of the SSEs is not guaranteed. On such an application, the proposed approach is advantageous to the existing approaches discussed in this paper.
However, there is room for improvement both from theoretical and practical viewpoints. Further research must focus on addressing the limitations of the current work listed below.
First, the Pareto-optimality of the proposed algorithm is not guaranteed. A necessary and sufficient condition for the Pareto-optimality of the algorithm must be derived. The development of a restart strategy to satisfy such a sufficient condition is desired.
Second, a computationally efficient and algorithmic update rule to select the next policy (cf., Equation 10 or Equation 9) is required. Practically, such an update similar to the policy iteration in single-agent MDPs, is desired.
Finally, our approach requires the knowledge of the follower’s best response and transition probability as in the previous studies. A sample approximation of the algorithm by reinforcement learning is desired to widen the applicability of the proposed approach.
Appendix A Proof of Theorem 5.1
Proof.
First, letting $V_{B}$ be fixed, we prove that $(TV)_{A}\stackrel{{\scriptstyle s}}{{=}}V_{A}$ implies $V_{A}\stackrel{{\scriptstyle s}}{{=}}\max_{f\in\mathcal{W}_{A}}V_{A}^{fR_{B}(f,V_{B})}$.
@@
=$\end{multlined}r_{A}(s,f(s),R_{B}(s,f(s),V_{B}))+\gamma_{A}\mathbb{E}_{s_{1}}\Bigl{[}r_{A}(s_{1},f(s_{1}),R_{B}(s_{1},f(s_{1}),V_{B}))+\gamma_{A}\mathbb{E}_{s_{2}}\left[V_{A}(s_{2})\right]\Bigl{]}$ (12c)
@@
$=\mathbb{E}^{fR_{B}(f,V_{B})}\left[\sum_{t=0}^{n-1}\gamma_{A}^{t}\hat{r}_{A}(s_{t},a_{t},b_{t})\middle|s_{0}=s\right]+\gamma_{A}^{n}\mathbb{E}^{fR_{B}(f,V_{B})}\Bigl{[}V_{A}(s_{n})\Bigl{|}s_{0}=s\Bigl{]}.$ (12d)
In the limit for $n\to\infty$, the first term of the right-hand side of Equation 12d converges to $V_{A}^{fR_{B}(f,V_{B})}(s)$, while the second term converges to 0777Since $\hat{r}_{A}$ is bounded and $\gamma_{A}\in[0,1)$, the inside of the expectation of Equation 12d is always bounded, which enables the exchange of the extreme and the expectation. Therefore, the same exchanges are performed in Equation 16d and Equation 37e.. As the choice of $f$ is arbitrary, we obtain $V_{A}\succcurlyeq\sup_{f\in\mathcal{W}_{A}}V_{A}^{fR_{B}(f,V_{B})}$.
@@
$=r_{A}(s,f^{*}(s),R_{B}(s,f^{*}(s),V_{B}))+\gamma_{A}\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,f^{*}(s),R_{B}(s,f^{*}(s),V_{B}))}\left[V_{A}(s^{\prime})\right].$ (13d)
@@
for all $s\in\mathcal{S}$ and for all $g\in\mathcal{W}_{B}^{d}$. By recursively applying Equation 15 for $V_{B}$ on the right-hand side, we have for all $s\in\mathcal{S}$
@@
$\end{multlined}r_{A}(s,R_{A}(s,V),g(s))+\gamma_{B}\mathbb{E}_{s_{1}}\Bigl{[}r_{A}(s_{1},R_{A}(s_{1},V),g(s_{1}))+\gamma_{B}\mathbb{E}_{s_{2}}\left[V_{B}(s_{2})\right]\Bigl{]}$ (16c)
@@
$\geq\mathbb{E}^{R_{A}(V)g}\left[\sum_{t=0}^{n-1}\gamma_{B}^{t}\hat{r}_{B}(s_{t},a_{t},b_{t})\middle|s_{0}=s\right]+\gamma_{B}^{n}\mathbb{E}^{R_{A}(V)g}\Bigl{[}V_{B}(s_{n})\Bigl{|}s_{0}=s\Bigl{]}.$ (16d)
In the limit for $n\to\infty$, the first term of the right-hand side of Equation 16d converges to $V_{B}^{R_{A}(V)g}(s)$, and the second term converges to 0. Therefore, $V_{B}(s)\geq V_{B}^{R_{A}(V)g}(s)$ holds for all $s\in\mathcal{S}$. Since $g\in\mathcal{W}_{B}^{d}$ is arbitrary, it holds that

### v2

@@ 6 Pareto-Optimal Policy Iteration
SSE policies do not always exist in general-sum stochastic games. Therefore, it is desired that an algorithm converges to an SSE policy if it exists and to a policy that is “reasonable” provided an SSE policy does not exist. First, we introduce the notion of Pareto-optimality as the reasonable target. A policy that approximates a Pareto-optimal value function with an arbitrary precision always exists, and it admits an SSE policy if an SSE policy exists. Then, we derive a policy improvement theorem for SSGs under best-response followers. Based on the policy improvement theorem, we design an algorithm that monotonically improves the policy toward Pareto-optimal policies.
Notation
We introduce the notation for the dominance relation. Let $v,v^{\prime}\in\mathcal{F}_{\mathcal{S}}$. If $v(s)=v^{\prime}(s)$ for all $s\in\mathcal{S}$, we express $v\stackrel{{\scriptstyle s}}{{=}}v^{\prime}$. If $v(s)\geq v^{\prime}(s)$ for all $s\in\mathcal{S}$, then $v$ weakly dominates $v^{\prime}$ and $v\succcurlyeq v^{\prime}$. If $v\succcurlyeq v^{\prime}$ and $\not\stackrel{{\scriptstyle s}}{{=}}v^{\prime}, vv$ strictly dominates $v^{\prime}$ and $v\succ v^{\prime}$.
@@ 6.1 Pareto-Optimal Policy
There always exists $\mathcal{PV}\neq\emptyset$. Moreover, $\mathcal{PV}\subseteq\partial\mathcal{V}$. A PO policy exists if and only if $\mathcal{PV}\cap\mathcal{V}\neq\emptyset$. For any PO value function $v^{*}\in\mathcal{PV}$, there always exists a policy whose value function is arbitrarily close to $v^{*}$. Meanwhile, there is no stationary policy that realizes a value function better than a PO value function. Moreover, an arbitrary PO policy is an SSE policy if there exist SSE policies, as shown in Proposition˜6.2. Therefore, PO value functions are a reasonable alternative to the SE value function. The proof is presented in Appendix˜B.
@@
Figure 2: Pareto optimal value functions in the two-state value function space. Thick blue line: Pareto optimal value functions. Red shaded area (the upper right of the dot in the box): the area of value functions dominating the value function represented by the black dot, which is empty for PO value functions.
@@ 6.2 Policy Improvement Theorem
Theorem˜6.3 is an extension of the policy improvement theorem (Sutton and Barto, 2018) in single-agent MDPs. With this theorem, we can confirm if a policy $f$ is better than a baseline $f^{\prime}$ without computing the value function for $f$. The proof is shown in Appendix˜C. The key idea behind its proof is to express the follower’s best response in terms of the leader’s policy and then apply reasoning analogous to the policy improvement theorem for single-agent MDPs.
Theorem 6.3.
@@
$(a)\quad Q_{A}^{f^{\prime}{\dagger}}(\cdot,f)\succcurlyeq V_{A}^{f^{\prime}{\dagger}}\implies V_{A}^{f{\dagger}}\succcurlyeq Q_{A}^{f^{\prime}{\dagger}}(\cdot,f)\succcurlyeq V_{A}^{f^{\prime}{\dagger}};$
$(b)\quad Q_{A}^{f^{\prime}{\dagger}}(\cdot,f)\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f^{\prime}{\dagger}}\iff V_{A}^{f{\dagger}}\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f^{\prime}{\dagger}};$
$(c)\quad Q_{A}^{f^{\prime}{\dagger}}(\cdot,f)\succ V_{A}^{f^{\prime}{\dagger}}\iff V_{A}^{f{\dagger}}-V_{A}^{f^{\prime}{\dagger}}\succ\gamma_{A}\mathbb{E}_{s^{\prime}}\left[V_{A}^{f{\dagger}}(s^{\prime})-V_{A}^{f^{\prime}{\dagger}}(s^{\prime})\right].$
A necessary condition for PO policies is given below, which shows that no further performance improvement can be made by Theorem˜6.3 (a) from PO policies. This proof is in Appendix˜D.
Theorem 6.4 (Necessary condition for PO policies).
@@
The following result shows a sufficient condition for PO policies. This proof is presented in Appendix˜E.
Theorem 6.5 (Sufficient condition for PO policies).
@@
$(\mathrm{i})\quad Q_{A}^{f^{\prime}{\dagger}}(\cdot,f)\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f^{\prime}{\dagger}};$
$(\mathrm{ii})\quad\exists s\in\mathcal{S},Q_{A}^{f^{\prime}{\dagger}}(s,f)<V_{A}^{f^{\prime}{\dagger}}(s)-\dfrac{\gamma_{A}}{1-\gamma_{A}}\delta(f,f^{\prime}),$
where $\delta(f,f^{\prime}):=\max\limits_{s\in\mathcal{S}}\mathbb{E}_{s^{\prime}\sim p(\cdot|s,f(s),R_{B}^{*}(s,f))}\left[Q_{A}^{f^{\prime}{\dagger}}(s^{\prime},f)-V_{A}^{f^{\prime}{\dagger}}(s^{\prime})\right].$
When $\gamma_{A}=0$, Theorem˜6.5 implies that no further policy update by Theorem˜6.3 (a) with strict improvement is a sufficient condition for PO policies. Combined with Theorem˜6.4, $f^{\prime}$ is a PO policy if and only if there is no room for policy improvement when $\gamma_{A}=0$.
@@ 6.3 Policy Iteration for PO Policies
We propose an approach that monotonically improves policy toward PO policies and is guaranteed to converge.
PO value functions are not always unique. To guide the algorithm, we introduce scalarization. Given a stationary policy $f\in\mathcal{W}_{A}$, we maximize a Pareto-compliant scalarization $\mathcal{L}$ of the state value on $f$, namely, $\mathcal{L}[V_{A}^{f{\dagger}}]$, where the Pareto-compliant scalarization [^4] For example, a weighted sum scalarization $\mathcal{L}[v]=\sum_{s\in\mathcal{S}}\alpha_{s}v(s)$ is a Pareto-compliant scalarization for $\boldsymbol{\alpha}:=\{\alpha_{s}\in\mathbb{R}_{>0}\}_{s\in\mathcal{S}}$ such that $\sum_{s\in\mathcal{S}}\alpha_{s}=1$. If we choose $\alpha$ such that $\alpha_{s}=\rho(s)$ for all $s\in\mathcal{S}$, then $\mathcal{L}[v]$ coincides with the expected return under the initial state distribution $\rho$. $\mathcal{L}:\mathrm{cl}\mathcal{V}\to\mathbb{R}$ is a continuous function such that $v\succ v^{\prime}\implies\mathcal{L}[v]>\mathcal{L}[v^{\prime}]$. Then, the maximum of the scalarized value in $\mathrm{cl}\mathcal{V}$ is obtained by a PO value function, namely, $\operatornamewithlimits{argmax}_{v\in\mathrm{cl}\mathcal{V}}\mathcal{L}\left[v\right]\subseteq\mathcal{PV}$.
@@
Selecting the next policy $f_{t+1}$ from $\mathcal{W}_{\succcurlyeq}(f_{t})$ guarantees $\mathcal{L}[V_{A}^{f_{t+1}{\dagger}}]\geq\mathcal{L}[V_{A}^{f_{t}{\dagger}}]$ in light of Theorem˜6.3 (a). Ideally, [^5] In practice, $\operatornamewithlimits{argmax}$ is not necessary, and sometimes it does not exist or is not computable. An alternative approach is to select a policy $f_{t+1}$ from a set $\mathcal{W}_{1-\epsilon}(f_{t})$ of $\epsilon$-optimal policies in terms of $\mathcal{L}$, such as $\mathcal{W}_{1-\epsilon}(f_{t}):=\bigg\{f\in\mathcal{W}_{\succcurlyeq}(f_{t})\mid\mathcal{L}\left[Q_{A}^{f_{t}{\dagger}}(\cdot,f)\right]-\mathcal{L}\left[V_{A}^{f_{t}{\dagger}}\right]\\ \geq(1-\epsilon)\sup_{f\in\mathcal{W}_{\succcurlyeq}(f_{t})}\bigg(\mathcal{L}\left[Q_{A}^{f_{t}{\dagger}}(\cdot,f)\right]-\mathcal{L}\left[V_{A}^{f_{t}{\dagger}}\right]\bigg)\bigg\}$ (7) for some $\epsilon\in[0,1)$. Notably, $\mathcal{W}(f_{t})=\mathcal{W}_{1}(f_{t})$. Therefore, the update in Equation 8 is a special case of Equation 7. If it is still intractable, selecting $f_{t+1}\in\mathcal{W}_{\succcurlyeq}(f_{t})$ s.t. $Q_{A}^{f_{t}{\dagger}}(\cdot,f_{t+1})\succ V_{A}^{f_{t}{\dagger}}$ ensures that (a) and (b) in Theorem 6.6 hold. the proposed algorithm selects $f_{t+1}$ from the set of policies that maximizes the scalarized value, namely,
@@
Theorem˜6.6 guarantees that the proposed algorithm always converges and that the limit satisfies the necessary condition for PO policies (i.e., $v_{\infty}\in\partial\mathcal{V}$). The proof is shown in Appendix˜F.
@@
Let $f_{0}\in\mathcal{W}_{A}$ be an arbitrary initial policy and a policy sequence $\{f_{t}\in\mathcal{W}(f_{t-1})\}_{t=1}^{\infty}$ be obtained by Equation˜7. Let $\{v_{t}=V_{A}^{f_{t}{\dagger}}\}_{t=0}^{\infty}$ be the corresponding value functions. Then, the following statements hold:
  • (a)
@@
  • (c)
    $v_{\infty}\in\partial\mathcal{V}$ and $\min_{s\in\mathcal{S}}v(s)-v_{\infty}(s)\leq\gamma_{A}\left(\max_{s\in\mathcal{S}}v(s)-v_{\infty}(s)\right)$ for all $v\in\mathcal{V}$;
  • (d)
@@
As shown in the proof (Appendix˜F), the algorithm ceases to improve if and only if $Q_{A}^{f_{t}{\dagger}}(\cdot,f)\stackrel{{\scriptstyle s}}{{=}}V_{A}^{f_{t}{\dagger}}$ for all $f\in\mathcal{W}_{\succcurlyeq}(f_{t})$. The algorithm terminates when it holds and returns $f_{t}$. The entire proposed algorithm is shown in Algorithm˜1.
Algorithm 1 Pareto-Optimal Policy Iteration
@@
4:  Compute $\mathcal{W}_{\succcurlyeq}(f_{t})$ in Equation˜6
@@
8:   Compute $\mathcal{W}(f_{t})$ in Equation˜8 or Equation˜7
@@
The convergence guarantee and the monotone improvement of Algorithm˜1 (Statements (a) and (b) in Theorem˜6.6) show the advantages of the proposed approach. However, compared with the policy iteration approach in single-agent MDPs, where the convergence to the optimal policy is guaranteed, the fixed point of Algorithm˜1 is not assured to be PO value functions. Rather, Theorem˜6.6 (c) guarantees that the limit is at the boundary $\partial\mathcal{V}$, including all the PO value functions. In a special case of $\gamma_{A}=0$, the limit is guaranteed to be a PO value function (Theorem˜6.6 (d)). The visualization of (c) and (d) is in Figure˜3.
This condition, the leader’s discount rate $\gamma_{A}=0$, is a condition on the leader’s learning process, meaning Algorithm˜1 does not need any assumptions on the follower’s learning process to guarantee the convergence and the leader’s “reasonable” performance. This property allows us to apply this algorithm to various applications modeled by decentralized SSGs in which we are the leader. We give an example of such applications in Appendix˜J.
There is another algorithmic difference from the policy iteration for the single-agent MDP. In single-agent MDPs, a state-wise maximization of the action value function of $f_{t}$, $a_{s}\in\operatornamewithlimits{argmax}_{a\in\mathcal{A}}Q^{f_{t}}(s,a)$ is performed to construct the next (deterministic) policy. This is invalid in SSGs because such a policy does not necessarily satisfy the condition of Theorem˜6.3 (a) and may not improve the values. Moreover, the optimal policy is not necessarily a deterministic one. Because of these differences, Equation˜8 cannot be simplified to the state-wise maximization. This can be a limitation from a practical viewpoint when Equation˜8 (or Equation˜7) is intractable due to the large search space of $\mathcal{W}_{A}$. We propose a practical strategy of splitting the policy space to find the next policy efficiently, avoiding exhaustive search over the entire space $\mathcal{W}_{A}$. Its details are provided in Appendix˜G
Figure 3: Visualization of Theorem˜6.6 (c) and (d). For $v_{\infty}$, there must not exist a value function in the red-shaded cone representing the area of $v$ violating the inequality in (c). The upper and lower slopes of the cone are $1/\gamma_{A}$ and $\gamma_{A}$, respectively. $v_{\infty}$ can be on the non-Pareto boundary (grey solid line) when $\gamma_{A}>0$ (left), but $v_{\infty}$ is on $\mathcal{PV}$ when $\gamma_{A}=0$ (right) because the cone equals the area of dominating value functions.
@@ 7 Limitation and Future Work
This study proposed a novel algorithm with a convergence guarantee in two-player general-sum SSGs under best-response followers. We introduced the notion of the Pareto-optimal value function to target it even if there are no SSEs, and developed an algorithm that monotonically improves the policy toward the Pareto front. While existing methods have little guarantee under best-response followers, especially in games where the SSEs do not exist, our proposed approach can monotonically improve the leader’s performance and guarantees its limit to be Pareto optimal in state values when the leader is myopic. To the best of our knowledge, this is the first theoretical guarantee in general-sum myopic-leader SSGs.
However, there is room for improvement both from theoretical and practical viewpoints. Further research must focus on addressing the limitations of the current work listed below. First, the Pareto-optimality of the proposed algorithm is not guaranteed. A necessary and sufficient condition for the Pareto-optimality of the algorithm must be derived. The development of a restart strategy to satisfy such a sufficient condition is desired. Second, a computationally efficient update rule to select the next policy (cf. Equation˜8 or Equation˜7) is required. Practically, a state-wise update similar to the policy iteration in single-agent MDPs is desired. Finally, our approach requires knowledge of the follower’s best response and transition probability, as in the previous studies. A sample approximation of the algorithm by reinforcement learning is desired to widen the applicability of the proposed approach.
Appendix A Proof of Theorem˜5.1
Proof A.1.
First, we prove that $(TV)_{A}\stackrel{{\scriptstyle s}}{{=}}V_{A}$ implies $V_{A}\stackrel{{\scriptstyle s}}{{=}}\max_{f\in\mathcal{W}_{A}}V_{A}^{fR_{B}(f,V_{B})}$ for a fixed $V_{B}$.
@@
$\geq\cdots$
[…]
In the limit for $n\to\infty$, the first term of the right-hand side of Equation˜10f converges to $V_{A}^{fR_{B}(f,V_{B})}(s)$, while the second term converges to 0666Since $\hat{r}_{A}$ is bounded and $\gamma_{A}\in[0,1)$, the inside of the expectation of Equation 10f is always bounded, which enables the exchange of the extreme and the expectation. Therefore, the same exchanges are performed in Equation 14e and Equation 19e.. As the choice of $f$ is arbitrary, we obtain $V_{A}\succcurlyeq\sup_{f\in\mathcal{W}_{A}}V_{A}^{fR_{B}(f,V_{B})}$.
@@
[…]
@@
for all $s\in\mathcal{S}$ and for all $g\in\mathcal{W}_{B}^{d}$. By recursively applying Equation˜13 for $V_{B}$ on the right-hand side, we have for all $s\in\mathcal{S}$
@@
$\geq\cdots$
[…]
In the limit for $n\to\infty$, the first term of the right-hand side of Equation˜14e converges to $V_{B}^{R_{A}(V)g}(s)$, and the second term converges to 0. Therefore, $V_{B}(s)\geq V_{B}^{R_{A}(V)g}(s)$ holds for all $s\in\mathcal{S}$. Since $g\in\mathcal{W}_{B}^{d}$ is arbitrary, it holds that